OpenStudy (anonymous):

solve for the indicated variable 5x - 7y =35 ; for y

OpenStudy (anonymous):

we cannot find two variables from a single equation

OpenStudy (anonymous):

solve for the indicated variable y

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

we can find y interms of x 5x - 7y =35 5x - 35 = 7y y = (5x - 35)/7

OpenStudy (anonymous):

Solve for y: 5x - 7y =35 Subtract 5x. -7y = -5x - 35 Divide by -7. y = 5/7 + 5

OpenStudy (anonymous):

Sorry: y = 5/7x + 5

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

to simplify that would it be y = 1 2/5 +5 and then y = 6 2/5?